What is "x"?
$8.16/3 lb. = x dollars/1 oz.

Answer by Alan3354(69443)

$8.16/3 lb. = x dollars/1 oz.
The units have to match.
Change the 3 lb to 48 oz
$8.16/48 = x/1
x = 8.16/48
x = $0.17 per ounce